An experiment was carried out to determine the mass of Sodium oxide in a solution sample. The sample was dissolved in water to make 200 cm³ of Sodium hydroxide. This solution required 500 cm³of Sulphuric acid to just neutralize it. The concentration of Sulphuric acid was 0.1mol/dm³ (0.1M)


1
Expert's answer
2022-05-24T03:54:03-0400

Na2O + H2O => 2NaOH

2NaOH + H2SO4 => Na2SO4 + 2H2O


1. We find amount substance(n, mol) of Sulphuric acid:

n (H2SO4) = V * CM

n – amount substance of sulphuric acid(mol);

V – solution volume of sulphuric acid = 500 cm3 = 50 dm3 = 50 litres;

CM – concentration of sulphuric acid = 0,1 mol/l .

n = 50 * 0,1 = 5 mol.


2. We find amount substance of sodium hydroxide. According to the above reaction, 1mol H2SO4 reacts with 2mol NaOH. How many moles NaOH react with 5mol H2SO4 :

1 mol H2SO4 — 2 mol NaOH

5 mol H2SO4 — X mol NaOH

X= 5*2/1 = 10 mol NaOH


3.We find the mass of Sodium oxide (Na2O)

According to the first reaction, 1 mol (62 g) of Na2O produces 2 mol of NaOH:

62 gr Na2O — 2 mol NaOH

X gr Na2O — 10 mol NaOH

X = 62 * 10 / 2 = 310 gr

ANSWER: 310 gr sodium oxide (Na2O)